# Soft-QMIX: Integrating Maximum Entropy For Monotonic Value Function Factorization

### v2

@@ Abstract
Multi-agent reinforcement learning (MARL) tasks often utilize a centralized training with decentralized execution (CTDE) framework. QMIX is a successful CTDE method that learns a credit assignment function to derive local value functions from a global value function, defining a deterministic local policy. However, QMIX is hindered by its poor exploration strategy. While maximum entropy reinforcement learning (RL) promotes better exploration through stochastic policies, QMIX’s process of credit assignment conflicts with the maximum entropy objective and the decentralized execution requirement, making it unsuitable for maximum entropy RL. In this paper, we propose an enhancement to QMIX by incorporating an additional local Q-value learning method within the maximum entropy RL framework. Our approach constrains the local Q-value estimates to maintain the correct ordering of all actions. Due to the monotonicity of the QMIX value function, these updates ensure that locally optimal actions align with globally optimal actions. We theoretically prove the monotonic improvement and convergence of our method to an optimal solution. Experimentally, we validate our algorithm in matrix games, Multi-Agent Particle Environment and demonstrate state-of-the-art performance in SMAC-v2.  [^1] Our code: https://github.com/WentseChen/Soft-QMIX
Figure 1: Soft-QMIX combines the QMIX value decomposition mechanism with maximum entropy RL under the CTDE framework. The figure explains the algorithm design rationale. The left figure illustrates a naive approach to applying maximum entropy RL in the CTDE context, where blue text represents the desired objectives, and black text indicates corresponding constraints, It reveals that incorporating maximum entropy RL under the CTDE framework implicitly constrains the global value to be the sum of local values, significantly limiting the expressiveness of the critic network. The right figure depicts the Soft-QMIX improvement, where Soft-QMIX first employs QMIX’s value decomposition mechanism to obtain $Q_{i}$. Given the meaningful order of QMIX’s output $Q_{i}$, such as possessing properties like IGM, an order-preserving transformation converts $Q_{i}$ to $f_{i}(Q_{i})$. This transformation is trained by the loss function that minimize the difference between $f_{i}(Q_{i})$ and $Q_{tot}.$
@@ 1 Introduction
In the context of solving cooperative tasks using deep multi-agent reinforcement learning (MARL), a standard approach is centralized training with decentralized execution (CTDE) [5, 23, 15]. Decentralized execution implies that during the execution phase, the decision-making of each agent is independent and limited to conditioning on local observations. Centralized training, on the other hand, allows the algorithm to access joint actions and global states while training the value function. Compared to decentralized training, which treats other agents as part of the environment, centralized training helps alleviate the instability often encountered in MARL training [30].
In most cooperative tasks, the multi-agent system receives a single joint reward, and MARL algorithms learn the corresponding joint value function based on this reward. Consequently, there is a need to design a credit assignment mechanism [2] that decomposes the joint value function into several local value functions, guiding the updates of the actor networks through these local value functions. The challenge in credit assignment lies in reconstructing the joint distribution from the combination of multiple marginal distributions, a process that typically involves approximation errors [32]. Therefore, past research has primarily focused on exploring the impact of mechanism design on the model’s capacity [27, 31].
Beyond model capabilities, exploration ability is also a critical factor affecting the performance of MARL algorithms. Maximum entropy RL [14], known for encouraging exploration [7], smoothing the landscape of the objective [1], and enhancing robustness [4], has been widely adopted in single-agent RL. However, integrating MARL with maximum entropy RL is not straightforward. As illustrated in Figure 1, using maximum entropy RL within the framework of CTDE will significantly limit the model’s abilities in a manner akin to VDN [28], which enforces that the sum of local q-values must equal the global q-value.
In this paper, we propose a value-based maximum entropy MARL method compatible with the CTDE framework, named Soft-QMIX. We adopt the QMIX [23] mechanism to train a centralized critic network and utilize its credit assignment mechanism to obtain local Q-functions. We also incorporate maximum entropy RL, utilizing stochastic policies and entropy regularization to improve exploration and robustness. To guarantee monotonically increasing expected returns during training and convergence to the optimal policy, an order-preserving transformation on the original local Q-functions is used for the final policy. Through this design, Soft-QMIX can benefit from both maximum entropy RL and QMIX’s credit assignment mechanism, while still ensuring convergence to the optimal strategy.
Our main contributions can be summarized as follows:
  • We introduce Soft-QMIX, a value-based maximum entropy MARL algorithm that aligns with the CTDE framework. It employs an order-preserving transformation on local q-values, enabling it to derive meaningful order from credit assignment and leverage the exploratory benefits provided by maximum entropy RL.
  • Theoretically, we demonstrate that our approach can monotonically improve the expected q-value and guarantee convergence to the optimal policy.
@@
  • Empirically, we confirm the effectiveness of our algorithm through tests on matrix games, MPE [18] and SMAC-v2 [3], where our method attains state-of-the-art results.
2 Related Works
Multi-Agent RL. In the CTDE framework, MARL primarily involves two common algorithmic approaches: policy gradient methods [17, 34, 15, 5, 13, 37] and value function decomposition methods [28, 23, 31]. Policy gradient methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence. On the other hand, Value function decomposition addresses the credit assignment problem by decomposing the global value function into multiple local value functions. VDN [28] assumes that the global Q-function is the sum of local Q-functions. QMIX [23, 20] permits the mixer function to be any function with non-negative weights. QTRAN [27] optimizes an additional pair of inequality constraints to construct a loss function. QPLEX [31] use the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize. The expressive power of the critic networks in these algorithms increases sequentially, but experimental results show that QMIX has better performance [9]. Therefore, in this work, QMIX is used as the mechanism for credit assignment.
Maximum Entropy RL. Maximum entropy RL [14] is demonstrated to have advantages such as encouraging exploration [7] and increasing robustness [4] in single-agent RL scenarios. In maximum entropy MARL, most works utilize an actor-critic architecture and typically aim to minimize the KL divergence between the local actor and the centralized critic [8, 6, 22, 36]. However, although these methods employ QMIX to train the critic, they are only capable of minimizing the discrepancy in expectation, thus cannot guarantee the IGM condition, which is central to the QMIX algorithm. Our approach is value-based and ensures that the IGM condition is met while employing maximum entropy MARL. Moreover, our method introduces global information into the local policy during training through a hyper-network, which most previous methods either ignored [35] or only applied in centralized training and centralized execution scenarios [16]. In appendix H, we provide a detailed comparison of our work with related works.
3 Preliminary
@@ 3.1 Multi-Agent Reinforcement Learning
Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [19]. Formally, a Dec-POMDP is represented as a tuple $(\mathcal{A},S,U,T,r,O,G,\gamma)$, where $\mathcal{A}\equiv\{1,...,n\}$ is the set of $n$ agents. $S$, $U$, $O$, and $\gamma$ are state space, action space, observation space, and discount factor, respectively.
During each discrete time step $t$, each agent $i\in\mathcal{A}$ chooses an action $u^{i}\in U$, resulting in a collective joint action $\mathbf{u}\in\mathbf{U}\equiv U^{n}$. The function $r(s,\mathbf{u})$ defines the immediate reward for all agents when the collective action $\mathbf{u}$ is taken in the state $s$. $\mathcal{P}(s,\mathbf{u},s^{\prime}):S\times\mathbf{U}\times S\rightarrow[0,\infty)$ is the state-transition function, which defines the probability density of the succeeding state $s^{\prime}$ after taking action $\mathbf{u}$ in state $s$. In a Dec-POMDP, each agent receives only partial observations $o\in O$ according to the observation function $G(s,i):S\times\mathcal{A}\rightarrow O$. For simplicity, if a function depends on both $o^{i}$ and $s$, we will disregard $o^{i}$. Each agent uses a policy $\pi_{i}(u^{i}|o^{i})$ to produce its action $u^{i}$ from its local observation $o^{i}$. Note that the policy should be conditioned on the history $o^{i}_{1:t}$. For simplicity, we refer to $o^{i}_{1:t}$ as $o^{i}_{t}$ or $o^{i}$ in the whole paper. The joint policy is denoted as $\pi_{jt}$. $\rho_{\pi}(s_{t},\mathbf{u_{t}})$ is used to represent the state-action marginals of the trajectory distribution induced by a policy $\pi$.
3.2 Individual-Global Condition
In this section, we define two key concepts in credit assignment.
@@ 3.2 Individual-Global Condition
In QMIX, to ensure the validity of IGM, the algorithm constrains the expressive capability of $Q_{tot}$ to the following form: $\frac{\partial Q_{tot}}{\partial Q_{i}}\geq 0,\forall i\in\mathcal{A}.$ Restricting the capacity of $Q_{tot}$ using the above inequality results in $Q_{tot}$ being a monotonic function. The monotonicity constraint serves as a sufficient, but not necessary, condition for IGM. Consequently, the monotonicity constraint is more stringent than those required by IGM.
Definition 3.2.
Individual-Global-Independence (IGI) For joint policy $\pi_{jt}$, if there exist individual policies $\left[\pi_{i}\right]_{i=1}^{n}$ such that the following holds:
$\pi_{jt}(\mathbf{u_{t}}|s_{t})=\prod_{i=1}^{n}\pi_{i}(u^{i}_{t}|o^{i}_{t}).$ (2)
Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGI for $\pi_{jt}$ under $s_{t}$.
Definition 3.1 characterizes the properties we expect a deterministic policy to possess. Definition 3.2, on the other hand, describes the independence of decision-making among agents during decentralized execution. IGI and the IGO [36] condition are very similar; we show in Appendix B that IGI is a sufficient condition for IGO.
3.3 Naive Maximum Entropy MARL Algorithm
In this section, we informally introduce a naive MARL algorithm. We use it to illustrate the issue that naively applying credit assignment to the maximum entropy RL algorithm will lead to limited credit assignment capability.
Assume a value-based framework, where $Q_{i}$ denotes the local Q-function and the function for credit assignment is denoted as $f^{c}$. Hence, the global Q-function, $Q_{tot}$, can be expressed as $Q_{tot}(\mathbf{u},s)=f^{c}\big{(}Q_{1}(u^{i},o^{i}),...,Q_{n}(u^{n},o^{n}),s\big{)}$. Based on maximum entropy RL, the global policy is defined as $\pi_{jt}(\cdot|s)\propto\text{softmax}\big{(}Q_{tot}(\cdot,s)\big{)}$ and the local policy as $\pi_{i}(\cdot|o^{i})\propto\text{softmax}\big{(}f^{t}_{i}(Q_{i}(\cdot,o^{i}),s)\big{)}$, where $f^{t}_{i}$ is an arbitrary function. Due to the IGI condition, we have $f^{c}(Q_{1},...,Q_{n},s)=\sum_{i=1}^{n}f^{t}_{i}(Q_{i},s)$. When $f^{t}_{i}$ is an identity mapping, the aforementioned algorithm is equivalent to VDN [28]. When $f^{t}_{i}$ represents a linear transformation, the algorithm becomes analogous to the DOP, an actor-critic method [33]. In fact, $f^{t}_{i}$ can be any function of $u^{i}$ and $o^{i}$, but cannot be a function of $u^{j}$ or $o^{j}$, where $j\neq i$. This constraint limits the expressiveness of the critic network.
4 Soft-QMIX
In this paragraph, we first briefly introduce the motivation behind the Soft-QMIX algorithm. Then, we detail the process of the Soft-QMIX algorithm and show that the expected returns of the centralized training are monotonically increasing and converge to the optimum. Finally, we propose a practical algorithm.
4.1 Motivation
As shown in Figure 2, we divide the decision-making process into two stages (green and blue blocks). In the first stage, agents rank the Q-values for all actions. We employ the value decomposition mechanism from QMIX to obtain local Q-values but only utilize the order of their outputs, not the specific Q-function values. In the second stage, without changing the order, we assign specific Q-values corresponding to each action. When selecting actions in decentralized execution, only the maximum value is required, necessitating just the first stage for execution. The second stage serves primarily to obtain an accurate estimated q-value. During this phase, we can utilize global information, as it is only used during training.
4.2 Soft Policy Iteration
The objective function of Maximum Entropy RL can be represented in the following form:
$J(\pi_{jt})$ $=\mathbb{E}_{(s_{t},\mathbf{u}_{t})\sim\rho_{\pi_{jt}}}\big{[}r(s_{0},\mathbf{u}_{0})+\sum_{t=1}^{\infty}\gamma^{t}\big{(}r(s_{t},\mathbf{u}_{t})+\alpha H\left(\pi_{jt}(\cdot|s_{t})\big{)}\right)\big{]},$ (3)
where $H(\pi_{jt})$ denotes the entropy of the joint policy $\pi_{jt}$, and the temperature parameter $\alpha$ is used to control the degree of randomness in the policy.
In the policy evaluation step, we aim to approximate the objective function described in Equation 3 using $Q_{tot}:S\times\mathbf{U}\rightarrow\mathbb{R}$. The soft state value function $V_{tot}:S\rightarrow\mathbb{R}$ is defined as follows:
$V_{tot}(s)=\mathbb{E}_{\mathbf{u}\sim\pi_{jt}(\cdot|s)}[Q_{tot}(s,\mathbf{u})-\alpha\log\pi_{jt}(\mathbf{u}|s)].$ (4)
Repeatedly applying the Bellman operation to update $Q_{tot}$ results in the soft $Q_{tot}$ function for policy $\pi_{jt}$. The Bellman operation is defined as follows:
$\mathcal{T}^{\pi_{jt}}Q_{tot}(s,\mathbf{u})=r(s,\mathbf{u})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}}[V_{tot}(s^{\prime})].$ (5)
In reference to the QMIX, a monotonicity constraint is imposed on the capacity of the critic network. Specifically, $Q_{\text{tot}}$ can be formulated as follows:
$Q_{tot}(s,\mathbf{u})=\text{Mixer}\left(Q_{1}(o^{1},u^{1}),...,Q_{n}(o^{n},u^{n}),s\right).$ (6)
The network structure of the Mixer function is the same as QMIX’s, and is elaborated in Appendix G.
Due to the monotonicity constraint, there is moderate error in estimating of $Q_{tot}$. To enhance the precision of value estimation, we introduce the TD($\lambda$[24] method. The estimation error between the maximum entropy objective function and the Bellman equation objective can be expressed as follows:
$\delta_{t}^{V}$ $=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1}),$ (7)
$\epsilon_{t}^{td(\lambda)}$ $=\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}(s_{t},\mathbf{u_{t}})-Q_{tot}(s_{t},\mathbf{u_{t}})=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta^{V}_{t+l}.$
Figure 2: The overall network architecture. Functions $g_{i}$ and $f_{i}$ are order-preserving transformations, ensuring the input and output dimensions are identical and have the same argmax values. The output of the $f_{i}$ function, after undergoing a softmax operation, becomes the policy, from which action $u_{t}^{i}$ is sampled. For decentralized execution, only the green network is needed, and the maximum values of their output are selected.
The derivation and implementation details of Equation 7 are provided in Appendix D. When combining off-policy methods with TD($\lambda$), adopting $\lambda>0$ requires importance sampling corrections [10]. Following previous work [9], we opt for a small $\lambda$ and assume that $\mathcal{T}^{\pi_{jt}}_{\lambda}$ is equivalent to $\mathcal{T}^{\pi_{jt}}$ in expectation. For simpilicity, we will use $\mathcal{T}^{\pi_{jt}}$ instead of $\mathcal{T}^{\pi_{jt}}_{\lambda}$ in theoretical discussions.
The following lemma demonstrates that by iteratively applying the Bellman equation, convergence can be achieved towards the maximum entropy objective $Q_{\text{tot}}$ for $\pi_{jt}$.
Lemma 4.1 (Joint Soft Policy Evaluation).
Let $Q^{0}_{jt}:S\times\mathcal{U}\to\mathbb{R}$ be a bounded mapping where $|\mathcal{U}|<\infty$. Define the iterative process $Q^{k+1}_{tot}=\mathcal{T}^{\pi_{jt}}(Q^{k}_{tot})$. It follows that the sequence $\{Q^{k}_{tot}\}_{k=0}^{\infty}$ converges to the joint soft-Q-function for policy $\pi_{jt}$ as $k\rightarrow\infty$.
Proof. See Appendix C. ∎
In Section 4.1, it was mentioned that decision-making is divided into two stages. During training, the policy is defined as $\pi^{train}_{i}(\cdot|o^{i},s)=\text{softmax}(f^{\text{stage2}}_{i}(f^{\text{stage1}}_{i}(\cdot|o^{i}),s)/\alpha)$, where $f^{\text{stage1}}_{i}$ is a function dependent on local observations, responsible for the first stage of decision-making. $f^{\text{stage2}}_{i}$, on the other hand, is a non-linear order-preserving transformation, responsible for the second stage of decision-making. Although $\pi^{train}_{i}$ depends on the global state, which conflicts with the definition in Dec-POMDP, it remains a valid option because it is utilized solely during training. In the subsequent paragraph, $\pi$ will denote $\pi^{train}_{i}$. $f^{\text{stage2}}_{i}$ can be further decomposed into two order-preserving transformations, $g_{i}$ and $f_{i}$. The output of $g_{i}$ is $Q_{i}$, and $f_{i}$ further transforms $Q_{i}$ into the final policy. The form of order-preserving transformation (OPT) is as follows:
$OPT_{i}(x,s)=W_{{}_{(1\times d_{1})}}^{2}\sigma(W^{1}_{{}_{(d_{1}\times 1)}}x_{{}_{(1\times d_{2})}}+b^{1})+b^{2},$ (8)
where $d_{1}$ is the hidden dimension, $d_{2}$ is the dimension of actions, $b^{1}$ and $b^{2}$ are the bias, $\sigma$ is an activation function, and $W^{1}$ and $W^{2}$ are non-negative weights. The implementation of the order-preserving transformation is similar to QMIX’s and is elaborated in Appendix G, where its weights and biases are provided by a hypernetwork. This hypernetwork takes the global state $s$ as input, making $g_{i}$ and $f_{i}$ functions dependent on the global state. Figure 2 presents the overall framework, where Agents $1$ to Agents $n$ correspond to $f_{i}^{\text{stage1}}$, and $g_{i}$ and $f_{i}$ correspond to $f_{i}^{\text{stage2}}$.
Since $g_{i}$ is an order-preserving transformation, and $[Q_{i}]_{i=1}^{n}$ satisfy the IGM condition for $Q_{tot}$, it follows that $\underset{\mathbf{u}}{\text{argmax}}\,Q_{tot}(\mathbf{u},s)=\left(\underset{u^{1}}{\text{argmax}}\,f_{1}^{\text{stage1}}(u^{1}|o^{1}),\ldots,\underset{u^{n}}{\text{argmax}}\,f_{n}^{\text{stage1}}(u^{n}|o^{n})\right)$. Therefore, during execution, only $f_{i}^{\text{stage1}}$ is needed.
During policy improvement, we update the joint policy towards the exponential of the joint Q-function. We constrain $\pi\in\Pi$, where $\Pi$ is the family of functions that satisfy the IGI condition. In the updating process, the updated policy is projected onto $\Pi$. Specifically, the update formula is as follows:

### v3

# ME-IGM: Individual-Global-Max in Maximum Entropy Multi-Agent Reinforcement Learning

@@ Abstract
Multi-agent credit assignment is a fundamental challenge for cooperative multi-agent reinforcement learning (MARL), where a team of agents learn from shared reward signals. The Individual-Global-Max (IGM) condition is a widely used principle for multi-agent credit assignment, requiring that the joint action determined by individual Q-functions maximizes the global Q-value. Meanwhile, the principle of maximum entropy has been leveraged to enhance exploration in MARL. However, we identify a critical limitation in existing maximum entropy MARL methods: a misalignment arises between local policies and the joint policy that maximizes the global Q-value, leading to violations of the IGM condition. To address this misalignment, we propose an order-preserving transformation. Building on it, we introduce ME-IGM, a novel maximum entropy MARL algorithm compatible with any credit assignment mechanism that satisfies the IGM condition while enjoying the benefits of maximum entropy exploration. We empirically evaluate two variants of ME-IGM: ME-QMIX and ME-QPLEX, in non-monotonic matrix games, and demonstrate their state-of-the-art performance across 17 scenarios in SMAC-v2 and Overcooked.
@@ 1 Introduction
Figure 1: The figure illustrates the improvement of our approach compared to existing maximum entropy MARL methods. The left figure shows a straightforward approach to applying maximum entropy MARL in the CTDE context, where blue texts represent the desired objectives, and black texts indicate corresponding constraints. It reveals that existing maximum entropy MARL methods under the CTDE framework implicitly constrain the global Q-value to be the sum of local Q-values (as in VDN), significantly limiting the expressiveness of the critic network. The right figure depicts the improvements in ME-IGM. ME-IGM first applies any credit assignment mechanism that satisfies the IGM condition, such as QMIX and QPLEX, to obtain $Q_{i}$ for each agent $i$. Given the meaningful order of local Q-values (for different actions), an order-preserving transformation $f_{i}$ converts $Q_{i}$ to $f_{i}(Q_{i})$ as the policy logits. This transformation is optimized using a loss function that minimizes the expected difference between $\sum_{i}f_{i}(Q_{i})$ and $Q_{tot}$, which guarantees monotonic policy improvement in maximum entropy MARL.
Collaborative multi-agent tasks, where a team of agents works together to complete a task and receives joint rewards, are inherently challenging for reinforcement learning (RL) agents. This difficulty arises due to the absence of individual reward signals for each agent and the significantly higher exploration requirements compared to single-agent scenarios. The challenge is further exacerbated by the exponential growth of the joint action space as the number of agents increases. In this work, we propose leveraging maximum entropy RL to promote exploration and adopting the individual-global-max (IGM) credit assignment mechanism (Rashid et al. 2020) to effectively distribute joint rewards among the agents.
Maximum entropy RL (Levine 2018), recognized for promoting exploration (Haarnoja et al. 2018), smoothing the objective landscape (Ahmed et al. 2019), and improving robustness (Eysenbach and Levine 2021), has been extensively applied in single-agent RL. However, as illustrated in Figure 1, naively extending maximum entropy RL to multi-agent settings under the centralized training and decentralized execution (CTDE) paradigm leads to uniform credit assignment, which limits the ability of multi-agent reinforcement learning (MARL) agents to learn from joint rewards.
On the other hand, numerous studies have explored the development of effective credit assignment mechanisms. Among them, the IGM condition has demonstrated strong empirical performance by aligning the maximum local Q-values with the maximum global Q-value. Previous works have primarily applied the IGM condition in deterministic policy settings, where policies are derived using the argmax of local Q-values. However, in maximum entropy MARL methods, stochastic local policies are learned, and misalignment can occur between local policies and the maximum global Q-values. That is, selecting joint actions based on the highest logits of local policies does not necessarily lead to maximizing the global Q-value. This limitation highlights a critical challenge in achieving effective credit assignment within the maximum entropy MARL framework. This challenge, as we will demonstrate, is not unique to IGM-based methods but also represents a broader obstacle for actor-critic-based maximum entropy MARL algorithms, such as FOP (Zhang et al. 2021), hindering their ability to achieve optimal coordination.
Our main contributions are summarized as follows:
  • We are the first to identify and empirically demonstrate a critical issue of existing maximum entropy MARL methods: the misalignment between stochastic local policies and the maximum global Q-values.
  • We introduce ME-IGM, the first CTDE maximum entropy MARL method that fully complies with the IGM condition. The core of ME-IGM is an order-preserving transformation (OPT), which maps local Q-values to policy logits while preserving their relative order. This ensures that selecting actions with the highest logits collectively leads to a joint action that maximizes the global Q-value. OPT can be seamlessly integrated into mainstream value decomposition MARL methods for a maximum entropy extension to enhance exploration. Additionally, we propose a theoretically grounded and straightforward objective for training these transformation operators.
@@
  • We confirm the effectiveness of our algorithm through extensive evaluation on matrix games, SMAC-v2 (Ellis et al. 2022), and Overcooked (Carroll et al. 2019), where our method attains state-of-the-art results.
2 Background
@@ 2.1 Multi-Agent Reinforcement Learning
Cooperative MARL can be formalized as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek and Amato 2016). Formally, a Dec-POMDP is represented as a tuple $(\mathcal{A},S,U,T,r,O,G,\gamma)$, where $\mathcal{A}\equiv\{1,...,n\}$ is the set of $n$ agents. $S$, $U$, $O$, and $\gamma$ are state space, action space, observation space, and discount factor, respectively.
During each discrete time step $t$, each agent $i\in\mathcal{A}$ chooses an action $u^{i}\in U$, resulting in a collective joint action $\mathbf{u}\in\mathbf{U}\equiv U^{n}$. The function $r(s,\mathbf{u})$ defines the immediate reward for all agents when the collective action $\mathbf{u}$ is taken in the state $s$. $\mathcal{P}(s,\mathbf{u},s^{\prime}):S\times\mathbf{U}\times S\rightarrow[0,\infty)$ is the state-transition function, which defines the probability density of the succeeding state $s^{\prime}$ after taking action $\mathbf{u}$ in state $s$. In a Dec-POMDP, each agent receives only partial observations $o\in O$ according to the observation function $G(s,i):S\times\mathcal{A}\rightarrow O$. For simplicity, if a function depends on both $o^{i}$ and $s$, we will disregard $o^{i}$. Each agent uses a policy $\pi_{i}(u^{i}|o^{i})$ to produce its action $u^{i}$. Note that the policy should be conditioned on the observation history $o^{i}_{1:t}$. For simplicity, we refer to $o^{i}_{1:t}$ as $o^{i}_{t}$ or $o^{i}$ in the whole paper. The joint policy is denoted as $\pi_{jt}$. $\rho_{\pi}(s_{t},\mathbf{u_{t}})$ is used to represent the state-action marginals of the trajectory distribution induced by a policy $\pi$.
2.2 Individual-Global-Max Condition
Individual-Global-Max (IGM) (Rashid et al. 2020) is a commonly used credit assignment method in value-decomposition-based cooperative MARL and is defined as follows:
@@ 2.2 Individual-Global-Max Condition
2.3 Related Works
Multi-Agent Credit Assignment: CTDE MARL primarily consists of two algorithmic branches: actor-critic methods (Lowe et al. 2017; Yu et al. 2022; Lin et al. 2023; Foerster et al. 2018; Kuba et al. 2021; Zhong et al. 2023) and value decomposition methods (Sunehag et al. 2017; Rashid et al. 2020; Wang et al. 2020). Actor-critic methods first learn a centralized critic network, and then distill local policies using losses like the KL divergence between local policies and the softmax of the global Q-value. On the other hand, value decomposition methods address the credit assignment problem by decomposing the global value function into multiple local value functions while satisfying the IGM condition. In particular, VDN (Sunehag et al. 2017) assumes that the global Q-function is the sum of local Q-functions. QMIX (Rashid et al. 2020) allows the mixing function, which aggregates local Q-values into the global Q-value, to be any function with non-negative partial derivatives, i.e., $\frac{\partial Q_{tot}}{\partial Q_{i}}\geq 0,\forall i\in\mathcal{A}.$ QTRAN (Son et al. 2019) introduces an additional set of inequality constraints to construct its loss function. QPLEX (Wang et al. 2020) uses the dueling architecture to decompose the Q-function, achieving the same expressive power as QTRAN while being easier to optimize. Our method, ME-IGM, enhances the exploration capability of these MARL algorithms by leveraging the maximum entropy RL framework and is compatible with all the aforementioned value decomposition methods.
Maximum Entropy MARL: Maximum entropy RL is demonstrated to encourage exploration in single-agent RL (Haarnoja et al. 2018). Most maximum entropy MARL works utilize an actor-critic architecture, minimizing the KL divergence between the local actors and the softmax of the centralized critic (He et al. 2022; Guo and Wu 2022; Zhang et al. 2021). However, as demonstrated later, these methods cannot guarantee that the joint action (at each state) determined by local policies maximizes the global Q-value, thus violating the IGM condition. In this case, we propose a novel order preserving transformation to ensure that the IGM condition is met while employing maximum entropy MARL.
3 Misalignment Between Local Policies and the Maximum Global Q-Value
The Misalignment Problem of Maximum Entropy MARL: While the IGM condition defines the relationship between local Q-values and the global Q-value, what is often more important in practice is the alignment between local policies and the global Q-value. In other words, joint actions that maximize the global Q-value can be obtained by querying local policies. Previous work in maximum entropy MARL (Zhang et al. 2021; Guo and Wu 2022) often decompose the global Q-value into individual local Q-values to satisfy the IGM condition. These local Q-values are then used to guide the training of local stochastic policies. Compared with value decomposition methods such as QMIX, they train stochastic policy networks in addition to the Q-functions, which can potentially improve multi-agent exploration in complex environments. However, while the value decomposition ensures alignment between global and local Q-values via the IGM condition, there is typically no explicit constraints imposed between the local policies and the local Q-values during the policy training phase. This lack of coupling often leads to misalignment during testing, where the local policies fail to consistently select joint actions that maximize the global Q-value, defeating the purpose of enforcing the IGM condition. Consequently, the learned policies may struggle to achieve optimal global coordination during execution.
Figure 2: Illustration of misalignment between local policies and the maximum global Q-value. When naively combining the IGM condition with maximum entropy MARL, local policies often select suboptimal joint actions, leading to a lower Q-value, depicted by the blue curve. In contrast, the optimal joint action should achieve the global-max Q-value, shown as the orange curve.
Empirical Demonstration: To analyze the performance gap caused by the misalignment between local policies and the maximum global Q-value, we evaluate the difference between the maximum global Q-value and the global Q-value selected by local policies in the SMAC-v2 zerg 5vs5 scenario, using a state-of-the-art Maximum Entropy MARL algorithm (Zhang et al. 2021). Figure 2 shows the results, averaged over five random seeds with standard deviations included. We can observe that there is a significant Q-value gap, emphasizing the potential for improvement by developing algorithms that enable the selection of joint actions with the highest global Q-value through local stochastic policies.
4 Order-Preserving Transformation
Here, we propose a novel order-preserving transformation (OPT) to solve the aforementioned misalignment problem between local policies and the maximum global Q-value.
4.1 Formulation of OPT
OPT is used to map local Q-values to the logits of local policies while preserving their relative order. Formally, the OPT is defined as:
$\text{OPT}(x,s)=W^{2}\sigma(W^{1}x+b^{1})+b^{2}$ (2)
Here, $W^{1}_{{}_{(d_{1}\times 1)}}$, $W_{{}_{(1\times d_{1})}}^{2}$, $b^{1}_{{}_{(d_{1}\times 1)}}$, and $b^{2}_{{}_{(1\times 1)}}$ are generated by a hyper-network [^3] For details of the hyper-network, please refer to the Appendix. which takes the global state $s$ as input. Consider $x$ as a local Q-value of dimension $(1\times d_{2})$, then $W^{1}x+b^{1}$ is a set of hidden variables of size $(d_{1}\times d_{2})$. Notably, when all entries in $W^{1}$ are non-negative, the relative order of elements in $x$ is preserved in each of the $d_{1}$ hidden variables. Next, a nonlinear activation function $\sigma$, such as ELU (Clevert 2015), is applied, followed by another order-preserving linear transformation using $W^{2}$ and $b^{2}$. Consequently, the final output, $\text{OPT}(x,s)$, maintains both the dimension and the relative order of the entries in the input $x$.
$\text{OPT}(x,s)$ is used as the logits of the local policy corresponding to the local Q-value $x$. Through aforementioned order-preserving design, each agent can select its action corresponding to the highest local policy logit, which is also the maximum point of its local Q-function. The local Q functions can be learned with any value decomposition methods satisfying the IGM condition, to ensure that the selected local actions collectively compose the optimal joint action that maximize the global Q-function. Thus, the proposed OPT addresses the aforementioned misalignment issue commonly found in maximum entropy MARL.
4.2 Training Scheme of OPT
Figure 3: The overall pipeline of ME-IGM.
Here, we first formulate the ideal objective of Maximum Entropy MARL and its practical counterpart that admits an analytical-form solution. We then analyze the gap between these two objective functions, providing a theoretically sound objective design for Maximum Entropy MARL based on OPTs.
Maximum Entropy MARL trains the individual and global Q-functions as in value decomposition MARL methods, while training a policy network in a Maximum Entropy RL framework based on the global Q-values. Formally, the joint policy is trained as follows: